OpenStudy (anonymous):

reduce this algebraic fraction b^2-3b-10/(b-2)^2*b-2/b-5

OpenStudy (campbell_st):

well factor the numerator in the 1st fraction \[\frac{(b - 5)(b + 2)}{(b - 2)(b - 2)} \times \frac{(b - 2)}{(b - 5)}\] cancel the common factors

OpenStudy (campbell_st):

of if you like its \[\frac{(b - 5)\times(b - 2)\times(b + 2)}{(b - 5) \times (b - 2)\times (b - 2)}\] and cancel the common factors for the answer

OpenStudy (anonymous):

I'm so lost :c

OpenStudy (campbell_st):

ok... so common factors are things that are in both the numerator and denominator... what appears factors (b - 2) is a factor... so what factors are in the denominator and numerator, that are the same..?

OpenStudy (anonymous):

b-2 and b-5?

OpenStudy (campbell_st):

great so remove 1 of each from the numerator and denominator what's left..?

OpenStudy (anonymous):

b+2/b-2?

OpenStudy (campbell_st):

that's the answer

OpenStudy (anonymous):

Dude, you're a life saver! Thank you so much!

OpenStudy (campbell_st):

glad to help